A fruit drink is 15% white grape juice. Use a system of equations to find the amount of pure white grape juice and the amount of this fruit drink needed to make 16 gallons of a new drink that is 21% white grape juice.
Answer by lwsshak3(11628)

A fruit drink is 15% white grape juice. Use a system of equations to find the amount of pure white grape juice and the amount of this fruit drink needed to make 16 gallons of a new drink that is 21% white grape juice.
***
let x=amt of pure white grape juice needed
let y=amt of 15% fruit drink needed
x+y=16
y=16-x
16-x=amt of 15% fruit drink needed
100%x+15%(16-x)=21%*16
x+2.4-.15x=3.36
.85x=0.96
x=1.13
y=16-x=14.87
amt of pure white grape juice needed=1.13 gal
amt of 15% fruit drink needed=14.87 gal